OpenStudy (anonymous):

5x - 5y = 10 3x - 2y = 2 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points) can someone help me with these steps please?

OpenStudy (aum):

5x - 5y = 10 ---- (1) 3x - 2y = 2 ---- (2) Part A: Let us replace the second equation by the sum of the second equation and 1/5 x the first equation: (1) * 1/5 = x - y = 2 (2): 3x - 2y = 2 Add: 4x - 3y = 4 ---- (3) Part B: Solve equations (1) and (2) for x and y. Solve equations (1) and (3) for x and y. Prove they are the same.

OpenStudy (anonymous):

ok?

OpenStudy (aum):

What did you get for x and y after solving equations (1) and (2)?

OpenStudy (anonymous):

I don't know how to solve, can u teach me

OpenStudy (anonymous):

thank you so that is part a, so now I need to use -2/3 and -8/3 to solve b?

OpenStudy (aum):

I will solve the equations (1) and (2). You use the same method to solve (1) and (3). 5x - 5y = 10 ---- (1) 3x - 2y = 2 ---- (2) Let us eliminate x. To do that multiply (1) by 3 and (2) by 5 and subtract: (1) * 3 = 15x - 15y = 30 (2) * 5 = 15x - 10y = 10 subtract: 0 - 5y = 20 divide by -5: y = 20/(-5) = -4 Put y in (1) and solve for x: 5x - 5(-4) = 10 5x + 20 = 10 5x= 10 - 20 = -10 x = -2 Solutions are x = -2; y = -4

OpenStudy (aum):

Part A has already been answered in my very first reply. The equivalent system of equations we created were: 5x - 5y = 10 and 4x - 3y = 4 For Part B) you need to solve the two original equations given in the problem, namely, 5x - 5y = 10 and 3x - 2y = 2 I have done that above and got x = -2 and y = -4. There is another thing you have to do for part B): Solve 5x - 5y = 10 and 4x - 3y = 4 and prove you get the same solution: x = -2 and y = -4.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Can someone help me with part b. I solved some of it but now im stuck. I got 20x-20y=40 and 20x-15y=20. what do I do next?